The University of Montana ski team has thirteen entrants in a men's downhill ski event. The coach would like the first, second, and third places to go to the team members. In how many ways can the thirteen team entrants achieve first, second, and third places?
Answer #1

here number of  ways can the thirteen team entrants achieve first, second, and third places =N(13 choice for 1st place,12 choice for 2nd place and 11 choice for 3rd place)=13*12*11 =1716